Project

General

Profile

06252014 Minutes

Attending: Michael, Hayes, Minnesotta (Dom, Susan, Nick) , Control Room (Fernanda and Gavin)

Diblock 7 added back into Partition 1 - Run
Other diblocks were added this week.

Far Detector -

Hits/spill - Jump in rate at about 9 am and back down at 10 am (Lights on the detector)

Jun 24 - drop in rate and slices (Running detector warm.)

Large gap in data in onmon plots (Not reco) - NUMI data time stamps were wrong (Evan has more detail)

P3 - One FEB was fluctuating in and out of run. (P3 is using the P2 timing, restarting runs on P2 may have caused this)

NDOS -
Timing peak looks good.

ND -
Shifter are supposed to keep ND running at the request of ND commissioning. P3 and P2 will start running.
Nearline is not ready to run from the near detector cluster.
OnMon and EVD are in the control room.
NearDet raw data is not being transferred right now.

Shifter can make watchdog report this week.